# Emulating Novel Topological Phases with Ultracold Fermions in Optical Lattices

### v2

@@ Abstract
We propose a feasible scheme utilizing the Floquet engineering platform with ultracold atoms in optical lattices to emulate both pair-hopping processes and singlet/triplet pairing interactions, aiming to simulate and control novel topologically nontrivial phases. Our large-scale density matrix renormalization group (DMRG) simulations reveal three distinct topological regimes: (i) a Majorana-enabled spin-density-wave (MS) phase characterized by exponentially localized edge charges, non-local single-fermion edge correlations, a doubly degenerate entanglement spectrum; (ii) a $z$-axis polarized triplet superconducting (TS) phase manifesting fractionalized edge spin; and (iii) a novel x-directional triplet superconducting (XTS) phase exhibiting fractional edge spins along the $x$-direction and non-local single-particle edge correlations, merging characteristics of the edge states from both MS and TS phases. Therefore, our work establishes a versatile framework for exploring exotic quantum phases and non-Abelian edge states in number-conserving systems.
@@
Recently, more and more attention has been focused on one-dimensional systems with a number-conserving setting Fidkowski et al. (2011); Sau et al. (2011); Cheng and Tu (2011); Kraus et al. (2013a); Ortiz et al. (2014); Kells (2015); Lang and Büchler (2015a); Keselman and Berg (2015a); Iemini et al. (2015a, 2017); Wang et al. (2017); Chen et al. (2018); Liu et al. (2019a); Yin et al. (2019); Lapa and Levin (2020), as these systems do not rely on an external pairing field. Ref. Cheng and Tu (2011) introduces a minimal particle number-conserving model. In this model, interchain pair tunneling serves as a key factor for MZMs by coupling two chains of a ladder with interacting spinless fermions. Subsequently, this novel pairing interaction was proposed in cold atomic systems by suppressing single-particle hopping by introducing offsets in the optical lattice Kraus et al. (2013a). However, its experimental realization remains challenging due to excessive laser requirements.
@@
The bosonic counterpart of this fermionic-type pairing (FTP) interaction, referred to as synthetic pair hopping (SPH) interaction, was proposed in our previous work Lin et al. (2020) through Floquet engineering in a two-component boson system. Compared to the implementation scheme of the SPH interaction Lin et al. (2020), it is easy to notice that by introducing nearest-neighbor repulsive interactions in a two-component fermion system and using the same driving scheme, we can realize this FTP interaction. In addition, reviewing the process of realizing the SPH interaction reveals that the synthetic anti-pairing interaction ($\hat{a}^{{\dagger}}_{s}\hat{b}_{s}\hat{b}^{{\dagger}}_{s}\hat{a}_{s}$) is inevitably generated simultaneously. Thus, a driving scheme similar to that for the SPH interaction to realize the FTP hopping interaction will inevitably generate simultaneously the fermionic-type anti-pairing hopping interaction. Interestingly, this fermionic-type anti-pairing hopping interaction can be expressed as a combination of spin-singlet pairing and $z$ component of spin-triplet pairing interaction (see a derivation in the Supplemental Material).
@@
On the other hand, searching for new topological phases and novel topological edge states is a fascinating research problem. Our previous analytical study Chen et al. (2018) revealed two distinct types of topological phases, i.e., $z$-directional triplet pairing superconducting (TS) phase and Majorana-spin-density-wave (MS) phase, in an interacting fermionic two-leg ladder system with FTP interaction, spin-singlet pairing (SSP), and spin-triplet pairing (STP) interactions. Under the mean-filed approximation, the TS phase (MS phase with half filling) is characterized by a fractional edge spin (charge) and can be interpreted as two Kitaev (Su–Schrieffer–Heeger ) chains Chen et al. (2018). A natural question is whether a new topological phase could emerge in this system that simultaneously exhibits the properties of these two states.
@@
In this Letter, we propose a Floquet engineering scheme in a one-dimensional spin-fermion system to emulate the FTP, both SSP and STP interactions, aiming to simulate and control novel topologically nontrivial phases in optical lattices. Three topologically nontrivial phases are revealed: the MS phase, the TS phase, and the $x$-directional triplet pairing superconducting (XTS) phase. Intriguingly, the XTS phase displays characteristics of both the MS and TS phases. Specifically, the XTS phase exhibits fractional edge spins along the $x$-direction and non-local single-particle edge correlations. Furthermore, various topologically trivial phases are also exposed.
@@
The effective Hamiltonian under high-frequency approximation can be simplified as (see a detailed derivation in the Supplemental Material SM )
@@
Apart from the above three phases, DMRG calculations give a richer phase diagram shown in Fig. 2. For $W<0$, we find a novel topological XTS phase that features fractional edge spin along $x$-direction and non-local single-particle edge correlations. The topological XTS, TS, and MS phases are protected by time-reversal symmetry (TRS) and spin parity symmetry ($P_{s}$) Liu et al. (2019b), each exhibiting twofold degenerate ground states under OBC. For $W>0$, we identify a normal superconducting (NSC) phase with coexisting triplet and singlet pairing, as well as a ferromagnetic superconducting (XFM) phase with spin polarization along the $x$-direction. Further details of the NSC, XFM, and SS phases are provided in the Supplemental Material SM .
@@
Topological MS phase.— To characterize distinct quantum phases, we measure charge densities, entanglement spectrumLaflorencie (2016); Pollmann et al. (2010), spin-spin correlations, single-particle and superconducting pairing correlations. As shown in Fig. 3(a), edge charge accumulation in the MS phase follows an exponential density profile. The topological nature of the MS phase also manifests itself in the doubly degenerate entanglement spectrum (ES) and non-local single-fermion correlations $G^{\sigma}_{1j}$ between the edges. The eigenvalues $\lambda_{\alpha}$ ($\alpha=1,...,\chi$) of the reduced density matrix $\rho_{L/2}$ with dimension $\chi$ yield the ES as $-ln\lambda_{\alpha}$. In Fig. 3(d), $G^{\sigma}_{1j}=\langle c^{\dagger}_{1\sigma}c_{j\sigma}\rangle$ ($j\in[1,L],\sigma=\uparrow,\downarrow$) attains considerable values around marginal sites while decays exponentially into the bulk, indicating localized Majorana edge states. Additionally, the ground state doubly degenerates due to selecting different local fermionic parities Iemini et al. (2015b); Lang and Büchler (2015b).
@@
Unlike higher-dimensional systems with long-range order, 1D systems exhibit quasi-long-range order (QLRO), characterized by power-law decaying order parameter correlations. The various superconducting pairing correlations are defined as follows: the spin-singlet pairing $\Phi_{S}(r)=\langle\Delta^{\dagger}_{S}(i)\Delta_{S}(j)\rangle$ with $\Delta_{S}(j)=\frac{1}{\sqrt{2}}(c_{j\uparrow}c_{j+1\downarrow}-c_{j\downarrow}c_{j+1\uparrow})$, and the triplet pairing $\Phi_{T,s}(r)=\langle\Delta^{\dagger}_{T,s}(i)\Delta_{T,s}(j)\rangle$ with $\Delta_{T,1}(j)=c_{j\uparrow}c_{j+1\uparrow}$, $\Delta_{T,0}(j)=\frac{1}{\sqrt{2}}(c_{j\uparrow}c_{j+1\downarrow}+c_{j\downarrow}c_{j+1\uparrow})$, $\Delta_{T,-1}(j)=c_{j\downarrow}c_{j+1\downarrow}$, where $r=|j-i|$ is the distance between site $i$ and site $j$. In the MS phase, $\Phi_{T,\pm 1}(r)$ decay in power law $\sim r^{-K_{SC}}$ with the superconducting exponent $K_{SC}\simeq 1.33$ in Fig. 3(b). $K_{SC}<2$ implies a divergent superconducting susceptibility at zero-temperature limit Arrigoni et al. (2004); Qu et al. (2022). Furthermore, the single-particle correlations are suppressed in the bulk, showing exponential decay. This indicates a nonzero gap for bulk single-fermion excitations, which guarantees the stability of the zero-energy edge modes. Fourier transforms of the spin-spin correlations $F(r)=\langle\boldsymbol{S}_{i}\cdot\boldsymbol{S}_{j}\rangle$ show a $2k_{F}$ spin-density wave (SDW) with Fermi momentum $k_{F}=\pi\rho/2$ SM . This feature is distinct from the previously discussed topologically $p$-wave superconducting phase induced by pair hoppingKraus et al. (2013b); Iemini et al. (2015b); Lang and Büchler (2015b).
@@
Topological TS phase.— The predominant QLRO in the TS phase is characterized by zero spin-$z$ triplet pairing superconducting correlations $\Phi_{T,0}(r)$, with a power exponent $K_{SC}\simeq 0.92$, as depicted in Fig. 4. This phase coexists with a weak singlet superconductivity, exhibiting a $K_{SC}$ slightly larger than 2. The single-particle correlations $G_{\sigma}(r)=\langle c^{\dagger}_{i\sigma}c_{j\sigma}\rangle$ show exponential decay behavior due to the absence of gapless fermionic mode in the bulk. At the system edges, the accumulation of $z$-component spin is observed with an opposite sign between the two degenerate ground states (GSs). Because there is an anomalous relation between time-reversal and local fermion parity operator Keselman and Berg (2015b). If the total number of fermions $N$ is even, total $\langle S^{z}\rangle=0$ and the integrated spin in each half of the system is close to half of an electron spin. If $N$ is odd, the two degenerate GSs with total $\langle S^{z}\rangle=\pm 1/2$ host an exact integrated spin $\langle S^{z}_{half}\rangle=\pm 1/4$ in each half. As illustrated in Fig. 4(a), the aforementioned edge spin polarization can induce SDW in bulk. To further substantiate the presence of a low-energy edge state, we add spin-$\downarrow$ fermion at site $i$ into each of the two GSs with an even particle number $N$. Subsequently, we compute the overlap with one of the GSs with a total particle number of $N+1$. The value of the overlap is nonzero at only one end depending on which GS is involved, clearly verifying a low-energy single-particle state localized at the edge (see Fig. 4(c)).
@@ Acknowledgements.
I Acknowledgements
This work is supported by the National Key Research and Development Program of China Grant No. 2022YFA1402204 and the National Natural Science Foundation of China Grant Nos. 12274086, 12004005, the Scientific Research Fund for Distinguished Young Scholars of the Education Department of Anhui Province No.2022AH020008, the Natural Science Foundation of Anhui Province under Grant Nos. 2008085QA26 and 2008085MA16, the University Synergy Innovation Program of Anhui Province under Grant No. GXXT-2022-039, and the open project of the state key laboratory of surface physics in Fudan University under Grant No. KF2021__\__08.

### v3

# Floquet-Engineered Hybrid Topological Orders with Majorana Edge Modes in Number-Conserving Fermionic Quantum Simulators

@@ Abstract
We develop an experimental protocol based on Floquet-engineered ultracold fermions in optical lattices, enabling the emulation of pair-hopping and competing singlet/triplet pairing interactions. Through large-scale density matrix renormalization group (DMRG) simulations, we uncover three emergent topological phases: (i) A Majorana-enabled spin-density-wave (MS) phase featuring exponentially localized edge charges, non-local fermionic edge correlations, and doubly degenerate entanglement spectra; (ii) A z-axis polarized triplet superconducting (TS) phase exhibiting fractionalized edge spins (S=1/4 per edge), two-fold ground state degeneracy and a bulk single-particle gap; (iii) A hybrid x-directional triplet superconducting (XTS) phase that uniquely combines fractional spin textures and Majorana-type edge correlations, defining a new universality class of hybrid orders in number-conserving systems. These findings establish a universal framework for engineering non-Abelian topological matter, crucially bypassing the need for external pairing fields while maintaining experimental feasibility with current cold-atom techniques.
@@
Recently, more and more attention has been focused on one-dimensional systems with a number-conserving setting Fidkowski et al. (2011); Sau et al. (2011); Cheng and Tu (2011); Kraus et al. (2013); Ortiz et al. (2014); Kells (2015); Lang and Büchler (2015); Keselman and Berg (2015); Iemini et al. (2015a, 2017); Wang et al. (2017); Chen et al. (2018); Liu et al. (2019a); Yin et al. (2019); Lapa and Levin (2020), as these systems do not rely on an external pairing field. Ref. Cheng and Tu (2011) introduces a minimal particle number-conserving model. In this model, interchain pair tunneling serves as a key factor for MZMs by coupling two chains of a ladder with interacting spinless fermions. Subsequently, this novel pairing interaction was proposed in cold atomic systems by suppressing single-particle hopping by introducing offsets in the optical lattice Kraus et al. (2013). However, its experimental realization remains challenging due to excessive laser requirements and requiring adiabatic eliminataion of the auxiliary degrees of freedom.
@@
The bosonic counterpart of this fermionic-type pairing (FTP) interaction, referred to as synthetic pair hopping (SPH) interaction, was proposed in our previous work Lin et al. (2020) through Floquet engineering with a continuous driving protocol in a two-component boson system. Subsequently, a pulse-sequence-based driving protocol Goldman et al. (2023) has been also theoretically shown to effectively implement the SPH interaction. Compared to the implementation scheme of the SPH interaction Lin et al. (2020), it is easy to notice that by introducing nearest-neighbor repulsive interactions in a two-component fermion system and using the same driving scheme, we can realize this FTP interaction. In addition, reviewing the process of realizing the SPH interaction reveals that the synthetic anti-pairing interaction ($\hat{a}^{{\dagger}}_{s}\hat{b}_{s}\hat{b}^{{\dagger}}_{s}\hat{a}_{s}$) is inevitably generated simultaneously. Thus, a driving scheme similar to that for the SPH interaction to realize the FTP hopping interaction will inevitably generate simultaneously the fermionic-type anti-pairing hopping interaction. Interestingly, this fermionic-type anti-pairing hopping interaction can be expressed as a combination of spin-singlet pairing and $z$ component of spin-triplet pairing interaction (see a derivation in the Supplementary Material).
@@
On the other hand, searching for new topological phases and novel topological edge states is a fascinating research problem. Our previous analytical study Liu et al. (2019a) revealed two distinct types of topological phases, i.e., $z$-directional triplet pairing superconducting (TS) phase and Majorana-spin-density-wave (MS) phase, in an interacting fermionic two-leg ladder system with FTP interaction, spin-singlet pairing (SSP), and spin-triplet pairing (STP) interactions. Under the mean-filed approximation, the TS phase (MS phase with half filling) is characterized by a fractional edge spin (charge) and can be interpreted as two Kitaev (Su–Schrieffer–Heeger ) chains Liu et al. (2019a). A natural question is whether a new topological phase could emerge in this system that simultaneously exhibits the properties of these two states.
@@
In this paper, we propose a Floquet engineering scheme in a one-dimensional spin-fermion system to emulate the FTP, both SSP and STP interactions, aiming to simulate and control novel topologically nontrivial phases in optical lattices. Three topologically nontrivial phases are revealed: the MS phase, the TS phase, and the $x$-directional triplet pairing superconducting (XTS) phase. Intriguingly, the XTS phase displays characteristics of both the MS and TS phases. Specifically, the XTS phase exhibits fractional edge spins along the $x$-direction and non-local single-particle edge correlations. Furthermore, various topologically trivial phases are also exposed.
@@
The effective Hamiltonian under high-frequency approximation can be simplified as (see a detailed derivation in the Supplementary Material SM )
@@
Apart from the above three phases, DMRG calculations give a richer phase diagram shown in Fig. 2. For $W<0$, we find a novel topological XTS phase that features fractional edge spin along $x$-direction and non-local single-particle edge correlations. The topological XTS, TS, and MS phases are protected by time-reversal symmetry (TRS) and spin parity symmetry ($P_{s}$) Liu et al. (2019b), each exhibiting twofold degenerate ground states under OBC. For $W>0$, we identify a normal superconducting (NSC) phase with coexisting triplet and singlet pairing, as well as a ferromagnetic superconducting (XFM) phase with spin polarization along the $x$-direction. Further details of the NSC, XFM, and SS phases are provided in the Supplementary Material SM .
@@
Topological MS phase.— To characterize distinct quantum phases, we measure charge densities, entanglement spectrumLaflorencie (2016); Pollmann et al. (2010), spin-spin correlations, single-particle and superconducting pairing correlations. As shown in Fig. 3(a), edge charge accumulation in the MS phase follows an exponential density profile. The topological nature of the MS phase also manifests itself in the doubly degenerate entanglement spectrum (ES) and non-local single-fermion correlations $G^{\sigma}_{1j}$ between the edges. The eigenvalues $\lambda_{\alpha}$ ($\alpha=1,...,\chi$) of the reduced density matrix $\rho_{L/2}$ with dimension $\chi$ yield the ES as $-ln\lambda_{\alpha}$. In Fig. 3(d), $G^{\sigma}_{1j}=\langle c^{\dagger}_{1\sigma}c_{j\sigma}\rangle$ ($j\in[1,L],\sigma=\uparrow,\downarrow$) attains considerable values around marginal sites while decays exponentially into the bulk, indicating localized Majorana edge states. Additionally, the ground state doubly degenerates due to selecting different local fermionic parities Iemini et al. (2015b); Lang and Büchler (2015).
@@
Unlike higher-dimensional systems with long-range order, 1D systems exhibit quasi-long-range order (QLRO), characterized by power-law decaying order parameter correlations. The various superconducting pairing correlations are defined as follows: the spin-singlet pairing $\Phi_{S}(r)=\langle\Delta^{\dagger}_{S}(i)\Delta_{S}(j)\rangle$ with $\Delta_{S}(j)=\frac{1}{\sqrt{2}}(c_{j\uparrow}c_{j+1\downarrow}-c_{j\downarrow}c_{j+1\uparrow})$, and the triplet pairing $\Phi_{T,s}(r)=\langle\Delta^{\dagger}_{T,s}(i)\Delta_{T,s}(j)\rangle$ with $\Delta_{T,1}(j)=c_{j\uparrow}c_{j+1\uparrow}$, $\Delta_{T,0}(j)=\frac{1}{\sqrt{2}}(c_{j\uparrow}c_{j+1\downarrow}+c_{j\downarrow}c_{j+1\uparrow})$, $\Delta_{T,-1}(j)=c_{j\downarrow}c_{j+1\downarrow}$, where $r=|j-i|$ is the distance between site $i$ and site $j$. In the MS phase, $\Phi_{T,\pm 1}(r)$ decay in power law $\sim r^{-K_{SC}}$ with the superconducting exponent $K_{SC}\simeq 1.33$ in Fig. 3(b). $K_{SC}<2$ implies a divergent superconducting susceptibility at zero-temperature limit Arrigoni et al. (2004); Qu et al. (2022). Furthermore, the single-particle correlations are suppressed in the bulk, showing exponential decay. This indicates a nonzero gap for bulk single-fermion excitations, which guarantees the stability of the zero-energy edge modes. Fourier transforms of the spin-spin correlations $F(r)=\langle\boldsymbol{S}_{i}\cdot\boldsymbol{S}_{j}\rangle$ show a $2k_{F}$ spin-density wave (SDW) with Fermi momentum $k_{F}=\pi\rho/2$ SM . This feature is distinct from the previously discussed topologically $p$-wave superconducting phase induced by pair hoppingLang and Büchler (2015); Kraus et al. (2013); Iemini et al. (2015b).
@@
Topological TS phase.— The predominant QLRO in the TS phase is characterized by zero spin-$z$ triplet pairing superconducting correlations $\Phi_{T,0}(r)$, with a power exponent $K_{SC}\simeq 0.92$, as depicted in Fig. 4. This phase coexists with a weak singlet superconductivity, exhibiting a $K_{SC}$ slightly larger than 2. The single-particle correlations $G_{\sigma}(r)=\langle c^{\dagger}_{i\sigma}c_{j\sigma}\rangle$ show exponential decay behavior due to the absence of gapless fermionic mode in the bulk. At the system edges, the accumulation of $z$-component spin is observed with an opposite sign between the two degenerate ground states (GSs). Because there is an anomalous relation between time-reversal and local fermion parity operator Keselman and Berg (2015). If the total number of fermions $N$ is even, total $\langle S^{z}\rangle=0$ and the integrated spin in each half of the system is close to half of an electron spin. If $N$ is odd, the two degenerate GSs with total $\langle S^{z}\rangle=\pm 1/2$ host an exact integrated spin $\langle S^{z}_{half}\rangle=\pm 1/4$ in each half. As illustrated in Fig. 4(a), the aforementioned edge spin polarization can induce SDW in bulk. To further substantiate the presence of a low-energy edge state, we add spin-$\downarrow$ fermion at site $i$ into each of the two GSs with an even particle number $N$. Subsequently, we compute the overlap with one of the GSs with a total particle number of $N+1$. The value of the overlap is nonzero at only one end depending on which GS is involved, clearly verifying a low-energy single-particle state localized at the edge (see Fig. 4(c)).
@@
Note added.— During the revision process, we noted the preprint Defossez et al. (2025), which presents a similar driving scheme for generating pair-hopping processes in a fermionic ladder system. Their numerical studies reveal a topological ground state featuring MZMs, reminiscent of our MS phase but lacking spin-density-wave order. Furthermore, a novel XTS phase has been discovered in our work, which uniquely supports both fractional edge spins and MZMs.
@@ Acknowledgements.
Acknowledgements.— This work is supported by the National Key Research and Development Program of China Grant No. 2022YFA1402204 and the National Natural Science Foundation of China Grant Nos. 12274086, 12004005, the Scientific Research Fund for Distinguished Young Scholars of the Education Department of Anhui Province No.2022AH020008, the Natural Science Foundation of Anhui Province under Grant Nos. 2008085QA26 and 2008085MA16, the University Synergy Innovation Program of Anhui Province under Grant No. GXXT-2022-039, and the open project of the state key laboratory of surface physics in Fudan University under Grant No. KF2021__\__08.